40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 20 21:43:25 2010, yuta, Update, CDS, checking whitening filters for MCs and messed up MC2ULSEN.pngMC2LRSEN.png
    Reply  Thu Oct 21 21:35:16 2010, yuta, Update, CDS, fixed whitening filter switches LLLR.pngContec23BO.png
    Reply  Fri Oct 22 00:59:01 2010, yuta, Update, CDS, new MEDM screens, new filter banks(whitening and dewhitening related) HappySideMEDM.png
       Reply  Fri Oct 22 23:29:48 2010, yuta, Update, CDS, fixed output filter switching 
          Reply  Mon Nov 1 23:31:05 2010, yuta, Update, CDS, checked input hardware filter in single frequency 
Message ID: 3748     Entry time: Wed Oct 20 21:43:25 2010     Reply to this: 3757   3758
Author: yuta 
Type: Update 
Category: CDS 
Subject: checking whitening filters for MCs and messed up 

(Joe, Yuta)

Background:
 We found that the damping servo for MC suspensions somehow worked when we turned off the 13Hz Chebyshev filters.
 But that does not meet our satisfaction, so we started checking every components.
 First of all is the whitening filters.
 If we turn on a digital whitening filter(WF), corresponding analog WF should turn off, and vice versa.

Reference:
 See DCC #D000210 for the analog circuit of WF. WF has 3Hz zero, 30Hz pole, 100Hz pole. MAX333A bypasses analog WF when supplied +15V(HIGH).

What we did:
 1. Compared the transfer function between MC2_SUSPOS_EXC and MC2_ULSEN_IN1 when digital WF on and off. When digital is on/off, analog should be off/on, so there should be difference but couldn't see.

 2. We went through the simulink model and found 2 mistakes in the logic. One is the conflict with other optics. Even if we turn on/off digital WF of MC2, it didn't switched analog WF of MC2. Two is the additional bit invert (but it turned out to be our misunderstanding).

 3. We (thought we) fixed it, rebuild it, and measured the TF again.(Attachment #1)

[Attached #1]
 The red/blue line is when digital WF is on/off. Blue should be bigger(+10dB @ 10Hz according to (analog) WTF) than red, but it was the opposite.

 4. To confirm that they are doing the opposite, I checked MAX333A input(pin#1,10,11,20) in "SUS PD Whitening Board" at 1X5-1-8B (which is for MC3) and found that switching is opposite. When I turned off/on the digital WF, MAX333A input was +15V/0V. It should be 0V/+15V.

 5. Also, I found that LLSEN digital WF switch switches LRSEN analog WF and vice versa.

[Attached #2]
 Transfer functions between MC2_SUSPOS_EXC and MC2_LRSEN_IN1 with;
   Red: LR digital off, LL digital on
   Blue: LR digital off, LL digital off
   Green: LR digital on, LL digital off
 As you can see, LL switch is the one which switches LR analog WF now.

Conclusion:
 Currently, digital WF on/off corresponds to analog WF on/off.
 Also, LL/LR digital WF switch is LR/LL analog WF switch now.


Next work:
 - fix the simulink models (or wiring)
 - check dewhitening filters

Schematic:
 - whitening
   MC1 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC2 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC3 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
      D000210 has switches for bypassing analog WT(3,100:3). HIGH to bypass.
 - dewhitening
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,3 UL/UR/LR/LL coils
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,2,3 SIDE coils
  (SimDW)(InvDW) ... -> LSC Anti-imaging Board(D000186) -> Universal Dewhitening Board(D000183) -> MC2 UL/UR/LR/LL coils
     D000316 has switches for bypassing 28Hz elliptic LPF. HIGH to bypass.
     D000186 is 7570Hz elliptic LPFs.
     D000183 has switches for bypassing dewhitening filter. HIGH to bypass.
 See this wiki page for more comprehensive setup.

Attachment 1: MC2ULSEN.png  43 kB  Uploaded Wed Oct 20 22:45:36 2010  | Hide | Hide all
MC2ULSEN.png
Attachment 2: MC2LRSEN.png  46 kB  Uploaded Wed Oct 20 22:45:57 2010  | Hide | Hide all
MC2LRSEN.png
ELOG V3.1.3-